OpenStudy (bj11684):

A computer downloaded 20 kilobytes of data in 4 seconds. If it downloads data at a constant speed, it can download ______ kilobytes of data in 15 seconds.

OpenStudy (anitas.):

what do you think? @bj11684

OpenStudy (czesc):

Set up a proportion. 20kilobytes/4seconds = ?kilobites/15seconds now solve for the question mark

OpenStudy (czesc):

20/4=X/15 5=X/15 15*5=X

OpenStudy (czesc):

x=75kilobytes, which is the answer

OpenStudy (anitas.):

no direct answers but you can lead to the answer.

OpenStudy (anitas.):

but great job :D